Find how many quarts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 40 quarts of 3% butterfat milk
Answer by wilft1(217)

two equations for this
x + y = 40
number of liters of each type that add up to 40 quarts
6x + 2y = 3
how much of each percentage type to make the 3 percent
start off by multiplying 3 by 40 and putting that answer in the bottom equation
x + y = 40
6x + 2y = 120
now, we are going to solve this using the addition method, we want to get rid of the y then solve for x, so lets multiply the top line by -2
-2x -2y = -80
6x + 2y = 120
add the two lines together
4x = 40
divide both sides by 4
x = 10
we now know the value of x, plug that number into our original equation and solve for y
x + y = 40
6(10) + 2y = 120
.....................
x + y = 40
60 + 2y = 120
..................
x + y = 40
2y = 60
.................
x + y = 40
y = 30
.............
x = 1 and y = 30
we now know that it will take 10 quarts of 6 percent and 30 quarts of 2 percent to make the 3 percent solution